OpenStudy (anonymous):

PLEASE HELP ME. FAN AND MEDAL The tires on Giuliana's bicycle have diameters of 24 inches. Giuliana is riding at a rate of 15ft per second. Approximately how many revolutions does her front wheel make each minute? Use 3.14 for pi. A. 286.6 B. 143.3 C. 11.9 D. 2

OpenStudy (anonymous):

@M4thM1nd

OpenStudy (anonymous):

@ganeshie8 @Luigi0210 @skullpatrol @iPwnBunnies @confluxepic @apoorvk @pooja195 @inkyvoyd @bibby @Mashy @mathmath333 @poopsiedoodle @jordanloveangel @HelpBlahBlahBlah @rational @linn99123 @Crissy15 @KyanTheDoodle @A_clan @NoelGreco @divu.mkr

OpenStudy (anonymous):

omg no one is answering its a 7th grade question

OpenStudy (anonymous):

is anyone gonna try to answer?

OpenStudy (anonymous):

as we know v=Rα

OpenStudy (anonymous):

and α=2πN where N is the number of revolution s per second

OpenStudy (anonymous):

then what do i do

OpenStudy (anonymous):

eegggh so amny marks

OpenStudy (anonymous):

so if we substitute we get v=(d/2)2πN

OpenStudy (anonymous):

whats with the infinite question marks

OpenStudy (anonymous):

replace and find N you get wait

OpenStudy (anonymous):

C 11.9 :)

OpenStudy (anonymous):

what do i replace with wat?

OpenStudy (anonymous):

25 with d?

OpenStudy (anonymous):

oh wait i got it thanks

OpenStudy (anonymous):

ok so i told you v=d2pπN so N=(V)/(d2π)

OpenStudy (anonymous):

and v=15 and d=24 and pie=3.14 so N=15/(24*3.14)=0.199rotations/s but they want it per minute so N=0.199*60=11.9rotations/min

OpenStudy (anonymous):

your welcome :)